# RECAP-KG: Mining Knowledge Graphs from Raw GP Notes for Remote COVID-19 Assessment in Primary Care

### v1

@@ Introduction
We now present the knowledge graph extraction phase of the pipeline (Figure 1 Stage 4), which extracts the knowledge graph from our internal Tree representation of the original passage. First, we propose a knowledge graph structure for GP notes. The nodes in the graph represent entities, and the edges represent relations between said entities. The information that we need to capture in the knowledge graph about a patient is what symptoms/signs the patient has, how severe these symptoms are, and for how long the patient has had them. The graph therefore contains four types of nodes: patient and symptom/sign entity nodes, symptom duration, and symptom severity.
  • The patient node corresponds to the subject of the medical summary who is assumed to always be the patient in question. There can only be one patient node in the graph. This node corresponds to the patient entity node in the parse tree.
  • Symptom/sign nodes represent any clinical condition/symptom attributed to the patient in the summary (e.g. shortness of breath, nausea, fever). There may be multiple signs/symptoms. These nodes correspond to the symptom/sign entity nodes in the parse tree. A symptom/sign node is always connected to the patient node via a ’has symptom’ or ’does not have symptom’ relation.
  • Symptom duration nodes are symptom descriptors that denote how long a corresponding symptom has been present (e.g. ’a week’, ’two days ago’). These nodes are typically generated from attachments. A symptom duration node is always connected to a symptom/sign entity node via a ’has duration’ relation.
  • Symptom severity nodes encode the severity and other properties of a corresponding symptom (e.g. ’slight’, ’severe’, ’intermittent’, ’continuous’). These nodes are typically generated from attachments. A symptom severity node is always connected to a symptom/sign entity node via a ’has severity’ relation.
@@
We now describe the process of extracting a knowledge graph from our internal parse tree representation. The knowledge graph extraction algorithm takes as input a list of augmented parse trees for the original text and individually extracts entity and relation instances from each tree. In order to do this, the algorithm heavily relies on open information extraction (OpenIE) - the process of extracting relations from text. Most importantly, OpenIE does not rely on a pre-defined schema in order to extract relations between two instances - instead, the relation name is simply the words linking the two. Just like constituency parsing, open information extraction performs better on a smaller input. It is a versatile language processing approach able to extract relations from any type of structured text. OpenIE can therefore be applied to both the complete passage and its smaller constituents, such as the core sentence and its attachments. Therefore, we will make use of the attachments generated in the parse tree augmentation stage in order to achieve the best performance. Namely, instead of extracting OpenIE relations from the whole original constituency parse tree, we will work with smaller segments, constructing sub-graphs. In our implementation, we use the AllenNLP OpenIE library(Gardner et al. 2018).
@@
The final output of the pipeline is a knowledge graph where the nodes are either entity nodes representing the patient and support phrases (signs/symptoms) or regular nodes corresponding to relevant words of the sentence parse tree. The edges record semantic relations between the nodes, such as a patient having a symptom and the symptom having a certain degree of severity.
@@
We want two answer two questions: (1) how well are the semantics of the original GP note preserved in the knowledge graph? and (2) is our bespoke KG extraction method as good as a language model for text understanding? We, therefore, evaluate the performance of our knowledge graph construction pipeline using two methods for the first and second questions respectively:
@@
    GP note comprehension: we translate the knowledge graphs into a symbolic representation, along with a set of predefined questions that a clinician might want to ask about a patient. We then run a symbolic solver on the result and compare the correctness of the answers with the performance of Bidirectional Encoder Representations from Transformers (BERT) (Zhang 2019) pre-trained for question answering. This method allows us to assess the text understanding of RECAP-KG by comparing it to one of the most widely used NLP models.
@@
We evaluate our model on CCAS - a dataset that contains structured and unstructured retrospective data on 315 000 COVID patients in the UK. The dataset includes GP consultation notes, as well as patient outcome (recovery, hospitalisation, death etc.). We use a subset of CCAS with 2004 entries for the evaluation which includes passages that contain at least one sentence that can be extended to the form ”Patient has ..’. Access to the CCAS data was obtained under approval from the Health Research Authority under section 251 of the NHS Act 2006. Privacy protection of the data was maintained by extracting text records from the CCAS EHR system, data then being linked per patient using NHS number to health data on hospital admissions and deaths (Hospital Episode Statistics) by NHS Digital. A hashing algorithm was used by Oxford university to Pseudonymise the data which was then made available for analysis in the ORCHID Secure Environment. Prior to making the data available, the Python routine Scrubadub (Beyond Accessed on 26 Februrary, 2023) was used to largely remove personal identifiers. HRA accepted that the remaining risk of disclosure was low and outweighed by the public benefit of using the data in the context of the COVID-19 pandemic.
@@
Two metrics are used to compute the similarity score between two texts: (a) Mean accuracy $\frac{1}{n}\sum_{i=1}^{n}sim(s_{1},s_{2})$, where $sim(s_{1},s_{2})$ is the cosine similarity between an original and deconstructed passage and (b) MSE computed as follows:
$\frac{1}{n}\sum_{i=1}^{n}{sim(s_{1},s_{2})}^{2}$
@@ Results
At present, on 2004 examples we achieve a 63% MSE and 79% mean cosine similarity when comparing sentences deconstructed from knowledge graphs and the original sentences.
@@
GP Notes Comprehension
In addition to evaluating preserved semantics of the original passage, we wish to evaluate whether our bespoke knowledge graph construction method is as expressive as a language model. Unlike our method, language models are not interpretable. We aim to assess the difference in performance between the two. We now describe the question-answering evaluation method where our model is compared to BERT(Zhang 2019) for question answering, trained on SQuAD(Rajpurkar et al. 2016) - a dataset containing over 100 000 questions for machine comprehension of text. In order to perform the evaluation, we have produced the questions that are related to the text that we would like our system to understand. The questions include eight yes/no questions and two multilabel questions that a healthcare professional might want to ask about a COVID-19 patient (Table 1).
Question Type
@@ 9 Does the patient have confusion? Y/N
10 How severe are the symptoms? Multilabel
11 When did the symptoms first start? Multilabel
@@
In order to perform question answering on our model, we translate the knowledge graphs, as well as question-answer ground truth pairs into a symbolic representation. The question-answer pairs are prompts for the language that has been extracted from the text. The formal definitions for this translation are detailed in Appendix Appendix. The translation produces a symbolic program where the question is encoded in zero or more background rules, the ground truth answer is encoded in a constraint, meaning that the program will fail if the ground truth is contradicted. The knowledge graph is encoded as a set of literals (symbolic facts). The program is then executed in the Answer Set Programming solver (Schaub Accessed on 26 Februrary, 2023), which produces a result that can either be ”satisfiable” - meaning that the facts extracted from the knowledge graph are consistent with the ground truth, or ”unsatisfiable” - meaning that there is a contradiction between the extracted facts and the ground truth constraint meaning that the information encoded in the knowledge graph is therefore incorrect. These steps are shown diagrammatically in Figure 7.
@@ Results
We evaluate the performance of the two models on two sets of yes/no and multilabel questions. The questions are grouped according to the answer type - in theory, the models should perform similarly on every question in a given group. The evaluation metric for the yes/no questions is accuracy, defined as
@@
For multilabel questions, the metric used is macro averaged precision, where precision is calculated for all classes individually and then averaged, defined as:
@@
We note from the results (Table 2) that our bespoke KG-construction method RECAP-KG consistently outperforms BERT for question answering. However, we do not achieve perfect accuracy, despite using a restricted subset of the data that our method has been curated for. This is due to the fact that our knowledge graph construction algorithm relies on good-quality parse trees, which are in turn reliant on good sentence structure. We enforce this sentence structure by splitting the original passage into chunks and thereafter extending each chunk to a full sentence by adding the subject and the verb if the verb is missing too (Figure 2 Steps 1.1, 1.2). However, our current sentence extension method is limited - it looks for a finite set of patterns in the constituency parse tree for the incomplete sentence only up to the first child of the root, under the assumption that the passage has been split into short chunks. When the assumption holds, the algorithm would find a match for one of the patterns, which is indeed the case for most entries in the data, as demonstrated by the high accuracy in the results. However, if not found, it may fail to add the subject or default to an incorrect verb form (e.g. ’patient is’ instead of ’patient has’). This is especially relevant for GP notes. Entries, where information about a patient is written in one long sentence with no conjunction/disjunction, are prevalent. We address this by splitting the passages at every delimiter. However, if no delimiters are present either, the passage containing one long sentence would not be split into chunks, passing a complex constituency parse tree to the sentence extension stage. The sentence extension stage might then fail to extend the sentence correctly, which would result in an incorrect internal tree representation.
@@
Table 2: Caption
@@
In this work, we have proposed a novel bespoke knowledge graph extraction method from GP notes and demonstrated that it outperforms BERT pre-trained for general question answering, while further offering a fully interpretable pipeline, unlike the black-box language models. Interpretability is a valuable property for healthcare, since it provides ease of validation when collaborating with domain experts and increases clinician trust. Therefore, RECAP-KG is a promising method for consolidating knowledge from GP notes. However, at present, our work possesses several limitations that need to be addressed. Firstly, the restricted nature of the data that our model is currently applied to (passages that contain at least one sentence that can be extended to the form ’Patient has ..’) does not reflect the full GP notes data, resulting in information loss due to sentences that do not fit this form not being processed. This can be addressed by extending the supported sentence structures, as in GP consultation notes they are typically one of a finite set (’Patient has to..’, ’Patient is..’, ’Patient’s X is Y’ etc.). Alternatively, the method can be improved by creating a sentence structure-agnostic pre-processing method. Secondly, we rely on AllenNLP’s(Gardner et al. 2018) constituency parser that uses SpaCy(spa Accessed on 5 June, 2019) parse trees - SpaCy relies on a neural network for sentence parsing, which inevitably creates some inconsistencies in creating parse trees (e.g. incorrect part-of-speech) and may thus affect the performance of our model. This can be addressed by adding additional constraints on the constituency parser.
@@
While our model demonstrates strong superiority over BERT for generic question answering trained on SQuAD(Rajpurkar et al. 2016) which contains over 100 000 questions for machine comprehension of text, this dataset is nevertheless not domain-specific. It would be of merit to compare RECAP-KG to a domain-specific version of BERT, such as BLUE BERT(Peng, Yan, and Lu 2019) that is pre-trained on PubMed abstracts and MIMIC-III (Johnson et al. 2016) clinical notes, after further pre-training it specifically for question answering. Finally, evaluating the performance on multiclass questions, such as ’What symptoms does the patient have?’ would provide more context on the text comprehension of RECAP-KG.
@@
The ultimate goal of our research is to produce knowledge graphs that can be used for providing more nuanced information about a patient to a clinical decision support system. Knowledge graphs extracted from datasets containing multiple patient entries can be used to infer new information about novel diseases, such as the novel coronavirus (COVID-19). Furthermore, automatic clinical decision system generation from knowledge graphs is an avenue we would like to explore. We have already developed a mechanism for translating RECAP-KG knowledge graphs into a symbolic representation. We envision building on this work and applying symbolic learning(Law, Russo, and Broda 2018) to a consolidation of facts extracted from the knowledge graphs in order to learn new insight about COVID-19 and to automatically generate clinical decision system support rules.
Appendix
Formal Definition of Graph and QA Translation to Symbolic Representation
We evaluate on two types of questions: (i) yes/no questions and (ii) multilabel questions.
  1. 1.
    Let RECAP-KG graph $G_{RECAP}$ be a tuple $(V,E)$, where $V$ is a set of $nodes$, and $E$ is a set of edges between the nodes $E\subseteq\{(s,t,d)\}$, s.t. $s,d\in V$ and $t\in\{hasSymptom,notHasSymptom,hasDuration$,
    $hasSeverity\}$ is the edge $type$.
  2. 2.
    Let a rule $r$ be
    $a_{1}\lor...\lor a_{n}\leftarrow b_{1},...,b_{m},\,not\,b_{m+1},...,not\,b_{r}$
    where $a_{1}...a_{n},\!b_{1},...,b_{m}$ are literals and $n,m,r\in\mathbb{N}$, $n,m\geq 0$, $r\geq m\geq 0$. A constraint is a rule without head literals. A set of rules is an ASP program(Law 2020).
  3. 3.
    Let a Question-Answer (qa) pair is a tuple $(q,a)$, where $q$ is a logic program and $a\in\{True,False\}$, which corresponds to the ground truth label.
  4. 4.
    Let $G$ be a graph s.t. $G=(V,E)$. Graph translation is a mapping $T_{G}:G\rightarrow P_{G}$ where $P_{G}$ is a conjunction of literals, such that for each edge $(s,t,d)\in E$:
    $T_{G}((s,t,d))$ $\rightarrow t(s,d).$
  5. 5.
    Let $QA$ be a set of yes/no question-answer pairs. QA translation is a mapping $T_{F}:QA\rightarrow P_{F}$ where $P_{F}\subseteq P$ and for each question-answer pair $(q,a)$:
[…]
We now provide additional definitions for multilabel questions. The two multilabel questions that we evaluate on are: ”When did the symptoms start?” (symptoms duration) and ”How severe are the symptoms?” (symptom severity).
Symptom Duration:
  1. 1.
    Aggregate atoms may appear in a rule body such that:
    $s_{1}\prec_{1}\alpha\{t_{1}:L_{1};...;t_{n}:Ln\}\prec_{2}s_{2},$
    where aggregate elements $t_{i}$ and $L_{i}$ are terms and literals respectively, $s_{1}$ and $s_{2}$ are terms, and $\alpha$ is a function. $\alpha$ is applied to the tuples $t_{i}$, given that $L_{i}$ is satisfied. $\prec_{1}$ and $\prec_{2}$ are comparison predicates to $s_{1}$ and $s_{2}$ respectively.
    NB: ${}^{\prime}s_{1}\prec_{1}^{\prime}$, $s_{1}\prec_{2}$, or both can be omitted - the comparison predicates then default to $\leq$
    Body aggregates act like positive literals when used together with the comparison predicate ’=’. Therefore, variable X is safe in
    $max(X)\,\texttt{:-}\,X=\#max\{t_{1}:L_{1};..;t_{n}:L_{n}\}.$
  2. 2.
    A symptom $S$ started $D$ days ago iff $has\_duration(S,D).$
  3. 3.
    The start of the first symptom is $d_{start}$ days ago iff
    $\forall s\in S,d\in\mathbb{N},0\leq d\leq 365,d_{start}=max_{d}(has\_duration(s,d))$
    where $S$ is the set of all possible symptoms.
  4. 4.
    Symptoms first started $D$ days ago iff:
    $-\displaystyle symptoms\_started(D)\texttt{:-}$-
    $D=\#max\{X:$ $has\_symptom(patient,S),$
    $has\_duration(S,X)\}.$
Therefore, given
  1. 1.
    QA pairs:
[…]
    $\forall_{i\in[1..n]}[(hasSymptom(s_{i}),True)]\},\forall_{i}[0\leq d_{i}]$
  2. 2.
    A graph $G=(V,E)$, where
    $\forall_{i\in[1..n]}[(patient,hasSymptom,s_{i}),(s_{i},hasDuration,d_{i})]\in E$
The ASP program
$-\displaystyle symptoms\_started(D)\texttt{:-}$-
$D$ =#$=\#max\{X:has\_symptom(patient,S),has\_duration(S,X)\}.$
$max\_duration(d_{max}).$
[…]
$has\_symptom(patient,s_{1}).$
$has\_symptom(patient,s_{n}).$
$has\_duration(s_{1},d_{1}).$
$has\_duration(s_{n},d_{n}).$
is satisfiable iff the grounding $symptoms\_started(d)$ is s.t. $d_{max}=d$.
Symptom Severity Given a QA pair $(hasSeverity(s),g).)$, and a graph $G=(V,E)$, where (s, hasSeverity, g_1) $\in E$, the ASP program
[…]
$has\_severity(s,g_{1}).$
is satisfiable iff $g_{1}=g$.
Therefore, the information encoded in a RECAP-KG graph is validated against a set of question-answer pairs by running the ASP program $P=P_{G}\cup P_{F}$. If $P$ is satisfiable, then the information encoded in the graph is correct with respect to the corresponding question.
ASP example
QA: {(hasSymptom(shortness_of_breath), True), (hasSymptom(fatigue), False)}
Graph: ({patient, shortness of breath, fatigue}, {(patient, fatigue, notHasSymptom), (patient, shortness of breath, hasSymptom)}
ASP program for validation:
:- not hasSymptom(patient, shortness_of_breath).
:- hasSymptom(patient, fatigue).
hasSymptom(patient, shortness_of_breath).
not hasSymptom(patient, fatigue).

### v2

@@ Introduction
We now present the knowledge graph extraction phase of the pipeline (Figure 1 Stage 4), which extracts the knowledge graph from our internal Tree representation of the original passage. First, we propose a knowledge graph structure for primary care physician notes. The nodes in the graph represent entities, and the edges represent relations between said entities. The information that we need to capture in the knowledge graph about a patient is what symptoms/signs the patient has, how severe these symptoms are, and for how long the patient has had them. The graph, therefore, contains four types of nodes: (I) the patient entity node that corresponds to the subject of the medical summary, (ii) symptom/sign entity nodes that represent any clinical condition/symptom attributed to the patient in the summary (e.g. shortness of breath, nausea, fever), (iii) symptom duration nodes that denote how long a corresponding symptom has been present (e.g. ’a week’, ’two days ago’), and (iv) symptom severity nodes that encode the severity and other properties of a corresponding symptom (e.g. ’slight’, ’severe’, ’intermittent’, ’continuous’).
@@
We now describe the process of extracting a knowledge graph from our internal parse tree representation. The knowledge graph extraction algorithm takes as input a list of augmented parse trees for the original text and individually extracts entity and relation instances from each tree. In order to do this, the algorithm heavily relies on open information extraction (OpenIE) - the process of extracting relations from text. Most importantly, OpenIE does not rely on a pre-defined schema in order to extract relations between two instances and can be applied to any kind of text. Therefore, it is applicable to both the complete passage and its smaller constituents, such as the core sentence and its attachments. Thus, instead of extracting OpenIE relations from the whole original constituency parse tree, we will work with smaller segments, constructing sub-graphs. In our implementation, we use the AllenNLP OpenIE library25.
@@
The final output of the pipeline is a knowledge graph where the nodes are either entity nodes representing the patient and support phrases (signs/symptoms) or regular nodes corresponding to relevant words of the sentence parse tree. The edges record semantic relations between the nodes, such as a patient having a symptom etc.
@@
We want two answer two questions: (1) how well are the semantics of the original primary care physician note preserved in the knowledge graph? and (2) is our bespoke KG extraction method as good as a language model for text understanding? We, therefore, evaluate the performance of our knowledge graph construction pipeline using two methods for the first and second questions respectively:
@@
    Consultation note comprehension: we translate the knowledge graphs into a symbolic representation, along with a set of predefined questions that a clinician might want to ask about a patient. We then run a symbolic solver on the result and compare the correctness of the answers with the performance of Bidirectional Encoder Representations from Transformers (BERT) pre-trained for question answering 12. This method allows us to assess the text understanding of RECAP-KG by comparing it to one of the most widely used NLP models.
@@
We evaluate our model on CCAS - a dataset that contains structured and unstructured retrospective data on 315 000 COVID patients in the UK. The dataset includes primary care physician consultation notes, as well as patient outcome (recovery, hospitalisation, death etc.). We use a subset of CCAS with 2004 entries for the evaluation which includes passages that contain at least one sentence that can be extended to the form ”Patient has ..’. Access to the CCAS data was obtained under approval from the Health Research Authority under section 251 of the NHS Act 2006. Privacy protection of the data was maintained by extracting text records from the CCAS EHR system, data then being linked per patient using NHS number to health data on hospital admissions and deaths (Hospital Episode Statistics) by NHS Digital. A hashing algorithm was used by Oxford university to Pseudonymise the data which was then made available for analysis in the ORCHID Secure Environment. Prior to making the data available, the Python routine Scrubadub 27 was used to largely remove personal identifiers. HRA accepted that the remaining risk of disclosure was low and outweighed by the public benefit of using the data in the context of the COVID-19 pandemic.
@@
We evaluate how close the reconstructed passage is to the original through semantic text similarity score (STS). The score is calculated by computing the cosine similarity between BERT28 embeddings for the original and reconstructed sentences. The STS scores are then averaged as follows: $\frac{1}{n}\sum_{i=1}^{n}STS(s_{1},s_{2})$, where $n$ is the number of passages in the original dataset.
@@ Results
At present, on 2004 examples we achieve a 79% mean STS score when comparing passages reconstructed from knowledge graphs and the original passages.
@@
Consultation Notes Comprehension
In addition to evaluating preserved semantics of the original passage, we wish to evaluate whether our bespoke knowledge graph construction method is as expressive as a language model. Unlike our method, language models are not interpretable. We aim to assess the difference in performance between the two. We now describe the question-answering evaluation method where our model will be compared to BERT28 for question answering, trained on SQuAD29 - a dataset containing over 100 000 questions for machine comprehension of text.
In order to perform the evaluation, we have produced the questions that are related to the text that we would like our system to understand. The questions include eight yes/no questions and two multiclass questions that a healthcare professional might want to ask about a COVID-19 patient (Table 1).
1 Question Type
@@ 9 Does the patient have confusion? Y/N
10 How severe are the symptoms? Multiclass
11 When did the symptoms first start? Multiclass
@@
In order to perform question answering on our model, we translate the knowledge graphs, as well as question-answer ground truth pairs into a symbolic representation. The question-answer pairs are prompts for the language that has been extracted from the text. The formal definitions are omitted from this paper due to their complexity. The translation produces a symbolic program where the question is encoded in zero or more background rules, the ground truth answer is encoded in a constraint, meaning that the program will fail if the ground truth is contradicted. The knowledge graph is encoded as a set of literals (symbolic facts). The program is then executed in the Answer Set Programming solver 30, which produces a result that can either be ”satisfiable” - meaning that the facts extracted from the knowledge graph are consistent with the ground truth or ”unsatisfiable” - meaning that there is a contradiction between the extracted facts and the ground truth constraint meaning that the information encoded in the knowledge graph is therefore incorrect (Figure 5).
@@ Results
We evaluate the performance of the two models on two sets of yes/no and multiclass questions. The questions are grouped according to the answer type - in theory, the models should perform similarly on every question in a given group. The evaluation metric for the yes/no questions is accuracy, defined as
@@
For multiclass questions, the metric used is macro averaged precision, where precision is calculated for all classes individually and then averaged, defined as:
@@
We note from the results (Table 2) that our bespoke KG-construction method RECAP-KG consistently outperforms BERT for question answering. However, we do not achieve perfect accuracy, despite using a restricted subset of the data that our method has been curated for. This is due to the fact that our knowledge graph construction algorithm relies on good-quality parse trees, which are in turn reliant on good sentence structure. We enforce this sentence structure by splitting the original passage into chunks and thereafter extending each chunk to a full sentence by adding the subject and the verb if the verb is missing too (Figure 2 Steps 1.1, 1.2). However, our current sentence extension method is limited - it looks for a finite set of patterns in the constituency parse tree for the incomplete sentence only up to the first child of the root, under the assumption that the passage has been split into short chunks. When the assumption holds, the algorithm would find a match for one of the patterns, which is indeed the case for most entries in the data, as demonstrated by the high accuracy in the results. However, if not found, it may fail to add the subject or default to an incorrect verb form (e.g. ’patient is’ instead of ’patient has’). This is especially relevant for notes in primary care. Entries, where information about a patient is written in one long sentence with no conjunction/disjunction, are prevalent. We address this by splitting the passages at every delimiter. However, if no delimiters are present either, the passage containing one long sentence would not be split into chunks, passing a complex constituency parse tree to the sentence extension stage. The sentence extension stage might then fail to extend the sentence correctly, which would result in an incorrect internal tree representation.
@@
Table 2: Comparison of RECAP-KG with BERT for question answering. There was no significant difference observed in performance between question within yes/no and multiclass question groups.
@@
In this work, we have proposed a novel bespoke knowledge graph extraction method from primary care consultation notes and demonstrated that it outperforms BERT pre-trained for general question answering, while further offering a fully interpretable pipeline, unlike the black-box language models. Interpretability is a valuable property for healthcare since it provides ease of validation when collaborating with domain experts and increases clinician trust. Therefore, RECAP-KG is a promising method for consolidating knowledge from primary care consultation notes. However, at present, our work possesses several limitations that need to be addressed. Firstly, the restricted nature of the data that our model is currently applied to (passages that contain at least one sentence that can be extended to the form ’Patient has ..’) does not reflect the full data, resulting in information loss due to sentences that do not fit this form not being processed. This can be addressed by extending the supported sentence structures, as in consultation notes they are typically one of a finite set (’Patient has to..’, ’Patient is..’, ’Patient’s X is Y’ etc.). Alternatively, the method can be improved by creating a sentence structure-agnostic pre-processing method. Secondly, we rely on AllenNLP’s25 constituency parser that uses SpaCy31 parse trees - SpaCy relies on a neural network for sentence parsing, which inevitably creates some inconsistencies in creating parse trees (e.g. incorrect part-of-speech) and may thus affect the performance of our model. This can be addressed by adding additional constraints on the constituency parser.
@@
While our model demonstrates strong superiority over BERT for question answering trained on SQuAD29, clinical versions of BERT are readily available. We have therefore fine-tuned BlueBERT (NCBI BERT) - a clinical version of BERT pre-trained on PubMed abstracts and MIMIC-III clinical notes32 for question answering. We have trained the model on the SQuAD dataset29 for 3 epochs with BertAdam optimiser with a learning rate 2e-05, train and evaluation batch size of 16. We intend to compare our method to this stronger benchmark as part of our future work. Finally, evaluating the performance on multilabel questions, such as ’What symptoms does the patient have?’ would provide more context on the text comprehension of RECAP-KG.
@@
The ultimate goal of our research is to produce knowledge graphs for providing more nuanced information about a patient to a CDSS. Knowledge graphs extracted from datasets containing multiple patient entries can be used to infer new information about diseases, e.g. novel coronavirus (COVID-19). Furthermore, automatic clinical decision system generation from knowledge graphs is an avenue we would like to explore. We have already developed a mechanism for translating RECAP-KG knowledge graphs into a symbolic representation. We envision building on this work and applying symbolic learning33 to a consolidation of facts extracted from the knowledge graphs to learn new insight about COVID-19 and to automatically generate CDSS support rules.